OpenStudy (anonymous):

If FI=6, and HI =12, calculate the length of GI

OpenStudy (anonymous):



OpenStudy (anonymous):

@Jack191 that is not one of my choices

OpenStudy (anonymous):

A.8 B.\[6\sqrt{2}\]

OpenStudy (anonymous):

C.12 D.10

OpenStudy (anonymous):

@Jack191 Feel what?

OpenStudy (anonymous):

feel the love

OpenStudy (anonymous):

well...the perpendicular to the base bisects the angle B. Consider triangle GHI. tanx=HI/GI . x=half of bisected angle B=90/2=45 degrees Just solve the equation tan45=HI/GI.